### v2

# Data-light Uncertainty Set Merging with Admissibility: Synthetics, Aggregation, and Test Inversion

Abstract
This article introduces a Synthetics, Aggregation, and Test inversion (SAT) approach for merging diverse and potentially dependent uncertainty sets into a single unified set. The procedure is data-light, relying only on initial sets and control levels, and it adapts to any user-specified initial uncertainty sets, accommodating potentially varying coverage levels. SAT is motivated by the challenge of integrating uncertainty sets when only the initial sets and their control levels are available—for example, when merging confidence sets from distributed sites under communication constraints or combining conformal prediction sets generated by different algorithms or data splits. To address this, SAT constructs and aggregates novel synthetic test statistics, and then derive merged sets through test inversion. Our method leverages the duality between set estimation and hypothesis testing, ensuring reliable coverage in dependent scenarios. A key theoretical contribution is a rigorous analysis of SAT’s properties, including a proof of its admissibility in the context of deterministic set merging. Both theoretical analyses and empirical results confirm the method’s finite-sample coverage validity and desirable set sizes.
Keywords: Admissibility, Conformal prediction, E-value, Finite-sample coverage, P-value, Synthetic test statistics.
1 Introduction
Uncertainty sets, such as confidence intervals and prediction intervals, are pivotal in statistical inference as they facilitate accurate representation and management of data variability. The integration of these sets has wide-ranging applications across various fields; however, considerable challenges emerge, especially when only the initial uncertainty sets and their control levels are available, along with possible intrinsic dependencies among the sets. To highlight the significance of set merging, we will first explore two prominent examples.
  1. (a).
    Distributed Learning with Communication Constraint. In distributed learning, the primary objective is to collaboratively make inferences using data distributed across different studies. Recent advancements have focused on distributed mean estimation (Cai and Wei,, 2024), prediction (Humbert et al.,, 2023), and causal inference (Xiong et al.,, 2023). A major challenge in distributed learning is the presence of communication constraints, which can arise from bandwidth limitations, privacy concerns, or cost considerations. These constraints restrict the amount of information that can be exchanged between studies. In certain scenarios, local sites can only transmit the confidence set and its associated confidence level to a central aggregator, highlighting the necessity for effective and data-light methods to merge these uncertainty sets for robust and reliable inferences.
  2. (b).
    Algorithmic Stability and Derandomization. Conformal prediction, pioneered by Vovk et al., (2005), has gained considerable popularity due to its minimal assumption requirements and its capability to provide finite-sample valid prediction sets for any black-box models. One of the most widely adopted versions is split conformal inference (e.g., Lei et al.,, 2018; Angelopoulos and Bates,, 2021), which is valued for its computational efficiency. However, the resulting prediction set can be influenced by the way the data is split. Additionally, variations in the prediction set may occur depending on the algorithm employed to compute the non-conformity score. To mitigate these issues, combining different prediction sets becomes a natural and necessary strategy.
This paper aims to develop an efficient and flexible method to combine $L$ different–potentially dependent–uncertainty sets into a single set, given only the initial uncertainty sets and their corresponding control levels. We term this method a data-light approach, emphasizing that it requires no access to raw data and the processes used to construct the initial uncertainty sets. Formally, let $Y$ denote the prediction target, and let $\mathcal{C}_{\ell,\alpha_{\ell}}$ represent any initial uncertainty set from the $\ell$-th study such that
$\mathbb{P}(Y\notin\mathcal{C}_{\ell,\alpha_{\ell}})\leq\alpha_{\ell},\quad\ell=1,\ldots,L,$ (1)
where $\alpha_{\ell}$’s are possibly varying control levels. The goal is to construct a merged set $\bar{\mathcal{C}}_{\alpha}$ such that $\mathbb{P}(Y\notin\bar{\mathcal{C}}_{\alpha})\leq\alpha$ for a pre-specified $\alpha\in(0,1)$, using only the available $\{(\mathcal{C}_{\ell,\alpha_{\ell}},\alpha_{\ell})\}_{\ell=1}^{L}$, while also ensuring that the merged set remains small in size.
1.1 Related Work
The aggregation of uncertainty sets has gained considerable interest recently, especially in the context of conformal prediction. Yang and Kuchibhotla, (2024) proposes selecting from nested conformal prediction sets the one with the smallest size, incorporating either coverage level adjustments or additional sample splitting. Liang et al., (2024) advances this approach by leveraging the properties of full conformal prediction, ensuring that coverage levels are always guaranteed with relatively small set sizes. Additionally, Stutz et al., (2021) focuses on training an optimal classifier that generates small conformal prediction sets by evaluating set size using a subset of mini-batch data during gradient descent. The main idea behind these methods is to choose an optimal non-conformity score to minimize the prediction set size, rather than merging the resulting sets. Furthermore, these approaches implicitly assume that all sets being aggregated have the same coverage guarantee.
In a different line of research, Chen et al., (2021); Bai et al., (2022); Fan et al., (2023); Kiyani et al., (2024) propose using constrained optimization approaches to directly minimize set size while maintaining coverage. However, these methods require access to the original data, which we do not assume to be available. Recently, Cherubin, (2019); Solari and Djordjilović, (2022); Gasparin and Ramdas, (2024) have explored merging uncertainty sets through majority voting. However, the admissibility of the majority voting method is not studied. In later sections, we show that our method can be viewed as a generalization of the voting method. In addition, we provide a theoretical analysis of the proposed approach, establishing key properties including its admissibility.
1.2 Our Method and Contributions
In this paper, we develop a novel data-light procedure for merging uncertainty sets, which we term Synthetics, Aggregation, and Test inversion (SAT). The SAT procedure operates in three core steps. First, we propose a novel class of synthetic test statistics that depend solely on the initial uncertainty sets and their associated coverage levels. These statistics are designed to mimic the true, underlying unknown statistics used to construct the initial sets, crucially without requiring access to the original data. Second, these synthetic test statistics are aggregated from multiple input sets based on appropriate aggregation functions. Finally, the merged uncertainty set is derived through the test inversion of the aggregated synthetic statistics.
This data-light approach requires minimal assumptions and only uses the initial uncertainty sets and their coverage levels, making it broadly applicable across various scientific domains. Our work offers several key contributions:
  1. (a).
    We propose a principled framework for merging uncertainty sets based on the duality between hypothesis testing and set estimation. This allows us to convert the set merging problem into the aggregation of test statistics, providing robustness to arbitrarily dependent input sets.
  2. (b).
    We introduce the concept of “Synthetic Statistics”, which bypasses the need for raw data by effectively mimicking unknown oracle test statistics using only the available initial sets and their coverage levels.
  3. (c).
    Importantly, by proposing the idea of synthetic statistics, we provide a theoretical analysis that proves the admissibility of the SAT procedure in a specific context, from which the admissibility of the majority voting method, as a special case of SAT, is derived. This represents a significant theoretical contribution, providing foundational support for a commonly used heuristic.
  4. (d).
    We establish a finite-sample coverage probability guarantee for the merged set derived from the SAT procedure, valid without model assumptions, and analyze the asymptotic size properties of the merged set.
1.3 Organization
The rest of the paper is organized as follows. Section 2 describes the proposed SAT procedure in detail. In Section 3, we establish the coverage guarantee and admissibility of SAT and analyze the size of the merged set theoretically. Sections 4 and 5 explore the numerical performance through simulations and the ImageNet__\__val dataset. We conclude the paper with a discussion of future directions in Section 6. Further extensions of the methods and theories, along with technical proofs and additional numerical results, can be found in the Supplementary Material.
1.4 Notations
Denote by $[n]=\{1,\dots,n\}$ for $n\in\mathbb{Z}^{+}$, and let $\mathbb{R}^{\geq 0}$ be the set of non-negative real numbers. Let $\operatorname{\mathds{1}}(\cdot)$ denote the indicator function. Let ${\bm{1}}_{L}$ be an $L$-dimensional vector of 1’s. For two positive sequences $\{a_{n}\}_{n\geq 1},\{b_{n}\}_{n\geq 1}$, write $a_{n}=O(b_{n})$ if there exists a constant $C>0$ such that $a_{n}/b_{n}\leq C$ for all $n$. Let $\|f\|_{L_{2}([0,1])}=\left\{\int_{0}^{1}f(x)^{2}{\rm d}x\right\}^{1/2}$ be the $L_{2}$-norm of $f$ over $[0,1]$, and denote $\|f\|_{L_{1}([0,1])}=\int_{0}^{1}|f(x)|{\rm d}x$ as the $L_{1}$-norm of $f$ over $[0,1]$. Denote by $\Phi(\cdot)$ and $z_{1-\alpha}$ the cumulative distribution function (CDF) and the $1-\alpha$ quantile of a standard normal random variable, respectively.
2 The SAT Procedure
This section proposes a SAT approach for merging uncertainty sets, utilizing only the initial sets and their corresponding control levels, while ensuring a guaranteed coverage probability. A broader class of uncertainty sets, where the notion of error extends beyond the miscoverage rate, will be further discussed in Section C.1 of the supplement.
Inspired by the duality between hypothesis testing and set estimation (Casella and Berger,, 2024), our proposed SAT procedure involves the following three steps:
Step 1 (Synthetics): Derive synthetic test statistics from initial uncertainty sets.
Step 2 (Aggregation): Aggregate test statistics from different studies.
Step 3 (Test Inversion): Merge uncertainty sets via test inversion of aggregated test statistics.
This sequence of steps outlines the intuition behind the SAT procedure, as detailed in Algorithm 1. Each step will be explained further in the subsequent subsections.
Algorithm 1 SAT Procedure
0:  The pairs $\{(\mathcal{C}_{\ell,\alpha_{\ell}},\alpha_{\ell})\}_{\ell\in[L]}$, candidate space $\mathcal{Y}$, a suitable aggregation function $G_{p}(\cdot)$ (or $G_{e}(\cdot)$), a control level $\alpha\in(0,1)$, an adjustment factor $\tau\in(0,1]$ (optional, $\tau$=1 by default).
0:  $\bar{\mathcal{C}}_{\alpha}\leftarrow\{\}$.
1:  for each candidate $y\in\mathcal{Y}$ do
2:     for each study $\ell\in[L]$ do
3:        Generate synthetic p-values $p_{\ell}(y)$ by (2) (or synthetic e-values $e_{\ell}(y)$ by (3)).
4:     end for
5:     Calculate $\bar{p}(y)=G_{p}\{\mathbf{p}(y)\}$ where $\mathbf{p}(y)=\{p_{1}(y),\ldots,p_{L}(y)\}$
                (or $\bar{e}(y)=G_{e}\{\mathbf{e}(y)\}$ where $\mathbf{e}(y)=\{e_{1}(y),\ldots,e_{L}(y)\}$).
6:     Update $\bar{\mathcal{C}}_{\alpha}\leftarrow\bar{\mathcal{C}}_{\alpha}\cup\{y\}$ if $\bar{p}(y)>\alpha$ (or if $\bar{e}(y)<\tau/\alpha$).
7:  end for
7:  Merged set $\bar{\mathcal{C}}_{\alpha}$.
2.1 Synthetic Statistics
A key component of the SAT procedure is the innovative construction of synthetic statistics, which translates initial uncertainty sets into test statistics through the duality of testing and interval estimation. Though the original data is not accessible, we can approximate the underlying (unknown) true statistics based on the initial sets and their levels, and construct synthetic versions of both p-values and e-values. In this section, we will detail this construction and examine the associated theoretical properties.
2.1.1 Synthetic p-value
A p-value $p\in[0,1]$ is a random variable that satisfies $\mathbb{P}(p\leq t)\leq t$ for all $t\in(0,1)$ under the null hypothesis [^2] With a slight abuse of terminology, the terms “p-values” and “e-values” below refer to both the random variables and their realized values, with the context clarifying the intended meaning.. It is well established that p-values can be employed to construct the uncertainty set, as formally stated in Propositions 7 and 8 in the subsequent sections. However, when only the uncertainty sets are provided, the true underlying statistics used to construct these sets, such as the p-values—referred to as “oracle p-values”—are not accessible. For each uncertainty set $\mathcal{C}_{\ell,\alpha_{\ell}}$ with control level $\alpha_{\ell}$, $\ell\in[L]$, we propose generating a random synthetic p-value that mimics such “oracle p-value” as follows:
$p_{\ell}(y):=p_{\ell}(y;\mathcal{C}_{\ell,\alpha_{\ell}},\alpha_{\ell})\sim{\rm Unif}\left(0,\alpha_{\ell}\right)\cdot\operatorname{\mathds{1}}(y\notin\mathcal{C}_{\ell,\alpha_{\ell}})+{\rm Unif}\left(\alpha_{\ell},1\right)\cdot\operatorname{\mathds{1}}(y\in\mathcal{C}_{\ell,\alpha_{\ell}}),\quad\forall y\in\mathcal{Y}.$ (2)
Intuitively, if $y\notin\mathcal{C}_{\ell,\alpha_{\ell}}$, it suggests that the underlying “oracle p-value” is likely small, so we generate the synthetic p-value from ${\rm Unif}\left(0,\alpha_{\ell}\right)$. On the other hand, if $y\in\mathcal{C}_{\ell,\alpha_{\ell}}$, it indicates that the “oracle p-value” is likely large, prompting us to generate the synthetic p-value from ${\rm Unif}\left(\alpha_{\ell},1\right)$. The following proposition demonstrates that the synthetic p-values constructed above satisfy the super-uniformity property.
Proposition 1.
Suppose (1) holds. Then, we have $\mathbb{P}\{p_{\ell}(Y)\leq t\}=t+\Delta_{\ell}(t)$ for all $t\in[0,1]$ and $\ell\in[L]$, where $p_{\ell}(\cdot)$ is the mapping defined in (2) and $\Delta_{\ell}(t)=\left\{1-\mathbb{P}\left(Y\notin\mathcal{C}_{\ell,\alpha_{\ell}}\right)/{\alpha_{\ell}}\right\}\cdot\{(t-\alpha_{\ell})\operatorname{\mathds{1}}\left(t>\alpha_{\ell}\right)-(t-t\alpha_{\ell})\}/(1-\alpha_{\ell})\leq 0.$
Remark 1.
As stated in the above proposition, exact uniformity ($\mathbb{P}\{p_{\ell}(Y)\leq t\}=t$) is achieved if the uncertainty set $\mathcal{C}_{\ell,\alpha_{\ell}}$ is exact, i.e., $\mathbb{P}(Y\notin\mathcal{C}_{\ell,\alpha_{\ell}})=\alpha_{\ell}$. More generally, Proposition 1 guarantees that $p_{\ell}(Y)$ is marginally super-uniform ($\mathbb{P}\{p_{\ell}(Y)\leq t\}\leq t$). However, it is important to note that this property may not hold for a fixed candidate value $y\in\mathcal{Y}$ when considering the sampling distribution.
Remark 2.
Given an uncertainty set derived from data, the synthetic p-values defined in (2) are randomly generated. The sensitivity of the merged set to this randomness is investigated in Section F of the supplement. A naive deterministic synthetic p-value can be constructed as $\alpha_{\ell}\cdot\operatorname{\mathds{1}}(y\notin\mathcal{C}_{\ell,\alpha_{\ell}})+\operatorname{\mathds{1}}(y\in\mathcal{C}_{\ell,\alpha_{\ell}})$. More sophisticated constructions of deterministic synthetic p-values, under certain assumptions, are detailed in Section B of the Supplement.
To compare the synthetic p-value with the “oracle p-value”, we use the following toy example to illustrate their differences in the context of merging confidence sets. Suppose we observe $X\sim\mathcal{N}(\theta^{*},1)$ and wish to construct a $(1-\alpha)$-level confidence interval for the mean parameter $\theta^{*}$. A natural choice of the “oracle p-value” for testing $H_{0}:\theta^{*}=\theta$ is $p^{\rm or}(\theta):=p^{\rm or}(\theta,X)=2\Phi(-|X-\theta|)$, and the corresponding confidence interval for $\theta^{*}$ can be written as $\mathcal{C}_{\alpha}=\{\theta:p^{\rm or}(\theta)>\alpha\}$. For the synthetic p-value generated by $p(\theta)\sim{\rm Unif}\left(0,\alpha\right)\cdot\operatorname{\mathds{1}}(\theta\notin\mathcal{C}_{\alpha})+{\rm Unif}\left(\alpha,1\right)\cdot\operatorname{\mathds{1}}(\theta\in\mathcal{C}_{\alpha})$, its comparison with $p^{\rm or}(\theta)$ is provided in Figure 1.
[…]
Figure 1: Comparison of oracle p-value (blue solid) and synthetic p-value (red dot). Left: null comparison with $\Delta(t)=F^{\rm or}(t)-F(t)$, where $F^{\rm or}(\cdot)$ and $F(\cdot)$ are CDFs of oracle and synthetic p-values, respectively. Middle & right: non-null comparisons with $\theta=1$ and $\theta^{*}=0$.
The left panel compares the CDFs of $p^{\rm or}(\theta)$ and $p(\theta)$ when $\theta=\theta^{*}$, showing that $p^{\rm or}(\theta)$ and $p(\theta)$ share the same uniform distribution in this case. The middle and right panels respectively compare the CDFs and density functions of $p^{\rm or}(\theta)$ and $p(\theta)$ when $\theta=1$ and $\theta^{*}=0$. We observe that $p^{\rm or}(\theta)$ is stochastically smaller than $p(\theta)$, indicating that $p^{\rm or}(\theta)$ is more powerful for testing $H_{0}:\theta^{*}=\theta$, while $p(\theta)$ offers a good approximation.
2.1.2 Synthetic e-value
Recently, e-value has gained popularity for hypothesis testing due to its favorable properties (Vovk and Wang,, 2021; Shafer,, 2021; Grünwald et al.,, 2020; Wang and Ramdas,, 2022). An e-value is a non-negative random variable with an expected value no greater than $1$ under the null hypothesis. Again, the true underlying e-value for each study, which we refer to as “oracle e-value”, is not available under our framework. Thus, for each uncertainty set $\mathcal{C}_{\ell,\alpha_{\ell}}$ with control level $\alpha_{\ell}$, $\ell\in[L]$, we propose to generate a synthetic e-value as follows:
$e_{\ell}(y):=e(y;\mathcal{C}_{\ell,\alpha_{\ell}},\alpha_{\ell})=\alpha_{\ell}^{-1}\cdot\operatorname{\mathds{1}}(y\notin\mathcal{C}_{\ell,\alpha_{\ell}}),\quad\forall y\in\mathcal{Y},$ (3)
where the corresponding function $e_{\ell}(\cdot)$ is referred to as an e-function. The expectation of $e_{\ell}(Y)$ is analyzed in the proposition below.
Proposition 2.
Suppose (1) holds. Then, we have $\mathbb{E}\{e_{\ell}(Y)\}\leq 1$ for all $\ell\in[L]$, where $e_{\ell}(\cdot)$ is the e-function defined in (3).
Remark 3.
The expectation in Proposition 2 is taken with respect to the randomness in both $Y$ and $\mathcal{C}_{\ell,\alpha_{\ell}}$. Notably, existing work primarily focuses on constructing e-values based on testing results with well-defined null hypotheses (e.g., Ren and Barber,, 2024; Bashari et al.,, 2024). In contrast, our framework differs considerably because establishing a well-defined null hypothesis in the current context is challenging due to the potential randomness of $Y$.
2.2 Aggregation of Synthetic Statistics
In this section, we provide a detailed discussion on how to aggregate synthetic p-values and e-values. The aggregated statistics will then be transformed back into a merged uncertainty set in the next step. Define $\mathbf{p}(Y)=\{p_{1}(Y),\dots,p_{L}(Y)\}\in[0,1]^{L}$ and $\mathbf{e}(Y)=\{e_{1}(Y),\dots,e_{L}(Y)\}\in[0,\infty)^{L}$, where $\mathbf{p}(\cdot)$ and $\mathbf{e}(\cdot)$ are mappings specified in Algorithm 1.
2.2.1 Aggregation under independence
We first consider independent case. Let $\mathbf{p}=(p_{1},\ldots,p_{L})$ be a vector of independent synthetic p-values. If each $p_{\ell}$ is a conventional p-value, then as pointed out in Vovk et al., (2022), we can define an aggregation function $G_{p}:[0,1]^{L}\mapsto[0,1]$ through the corresponding rejection regions. More precisely, given any increasing collection of Borel lower sets $\{R_{\alpha}\subseteq[0,1]^{L}:\alpha\in(0,1)\}$, if $\mathbb{P}(\mathbf{p}\in R_{\alpha})\leq\alpha$ for any $\alpha\in(0,1)$ under the null hypothesis, then $G_{p}(\mathbf{p})=\inf\{\alpha\in(0,1):\mathbf{p}\in R_{\alpha}\}$ defines a valid p-value aggregation function. Synthetic p-values can be aggregated using the same idea. Specifically, we consider the rejection regions of the following form:
$R_{\alpha}=\left\{\mathbf{p}\in[0,1]^{L}:\sum_{\ell=1}^{L}S_{\ell}(p_{\ell})\geq c_{1-\alpha}(\{S_{\ell}\}_{\ell\in[L]})\right\},$ (4)
where $S_{\ell}:[0,1]\mapsto\mathbb{R}$ is decreasing and $c_{1-\alpha}(\{S_{\ell}\}_{\ell\in[L]})={\rm Quantile}(1-\alpha;\sum_{\ell=1}^{L}S_{\ell}(U_{\ell}))$ with $U_{\ell}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}{\rm Unif}[0,1]$. Notably, (4) encompasses some of the most widely used aggregation methods for the conventional p-values. For example, by setting $S_{\ell}(t)=-2\log t$, we obtain Fisher’s aggregation function $G_{p}(\mathbf{p})=1-F_{\chi^{2}_{2L}}(-2\sum_{\ell=1}^{L}\log p_{\ell})$, where $F_{\chi^{2}_{2L}}$ denotes the CDF of a centered $\chi^{2}$-random variable with $2L$ degrees of freedom (Fisher,, 1948); by taking $S_{\ell}(t)=-\lambda_{\ell}\cdot\Phi^{-1}(t)$, where $\lambda_{\ell}$’s are some positive constants, we obtain the Lipták’s method $G_{p}(\mathbf{p})=\Phi\Big{\{}\sum_{\ell=1}^{L}\lambda_{\ell}\cdot\Phi^{-1}(p_{\ell})/\sqrt{\sum_{\ell=1}^{L}\lambda_{\ell}^{2}}\Big{\}}$ (Lipták,, 1958). For detailed comparisons of various aggregation methods under independence, see, for example, Heard and Rubin-Delanchy, (2018).
The following proposition shows that the aggregated synthetic p-value via (4) is still marginally super-uniform.
Proposition 3.
Suppose (1) holds. Let $\bar{p}(Y):=G_{p}\{\mathbf{p}(Y)\}=\inf\{\alpha\in(0,1):\mathbf{p}(Y)\in R_{\alpha}\}$, where $R_{\alpha}$ is defined in (4). If the entries of $\mathbf{p}(Y)$ are mutually independent, then $\mathbb{P}\{\bar{p}(Y)\leq t\}\leq t$ for all $t\in[0,1]$.
We next turn to the aggregation of independent synthetic e-values. The most popular method for aggregating independent conventional e-values is through multiplication or averaging (Vovk and Wang,, 2021). Similarly, synthetic e-values can be aggregated by
$\bar{e}_{k}:=G_{e}(\mathbf{e};k)=\binom{L}{k}^{-1}\sum_{\mathcal{I}_{k}\in\mathcal{B}_{k}}\prod_{\ell\in\mathcal{I}_{k}}e_{\ell},$ (5)
where $\mathbf{e}=(e_{1},\ldots,e_{L})$ is a vector of independent synthetic e-values and $\mathcal{B}_{k}$ is the set of all $k$-element subsets of $[L]$. We abbreviate $\bar{e}_{k}$ and $G_{e}(\cdot;k)$ as $\bar{e}$ and $G_{e}(\cdot)$, respectively, whenever $k$ is fixed and does not influence the calculations. The validity of this aggregation method is established by the following proposition.
Proposition 4.
Suppose (1) holds. Let $\bar{e}_{k}(Y)=G_{e}\{\mathbf{e}(Y);k\}$ be defined as in (5). If the entries of $\mathbf{e}(Y)$ are mutually independent, then $\mathbb{E}\{\bar{e}_{k}(Y)\}\leq 1$ for any pre-determined $k\in[L]$.
When $Y$ is a fixed but unknown parameter and each of the $L$ studies independently collects data to construct $\mathcal{C}_{\ell,\alpha_{\ell}}$, the entries in $\mathbf{p}(Y)$ and $\mathbf{e}(Y)$ are independent, making the aggregation methods discussed in this section appropriate. However, if $Y$ is random, the entries in $\mathbf{p}(Y)$ and $\mathbf{e}(Y)$ are generally not independent, even if each study independently collects data to construct $\mathcal{C}_{\ell,\alpha_{\ell}}$. This is because the randomness of $Y$ introduces dependence among the entries via their shared dependency on $Y$. Thus, it is crucial to account for this to ensure the validity of the aggregation method. The aggregation approach under dependence will be discussed next.
2.2.2 Aggregation under dependence
We now explore methods for aggregating synthetic statistics under arbitrary dependence. Note that, more sophisticated aggregation methods can be employed if specific dependence structures are assumed as discussed in Section A.2 of the supplement. Let $\mathbf{p}=(p_{1},\ldots,p_{L})$ be a vector of possibly dependent synthetic p-values. Consider a family of regions defined as follows:
$R_{\alpha}=\left\{\mathbf{p}\in[0,1]^{L}:\sum_{\ell=1}^{L}\lambda_{\ell}\cdot f_{\ell}\left(\frac{p_{\ell}}{\alpha}\right)\geq 1\right\},$ (6)
where $\lambda_{\ell}$’s are non-negative numbers satisfying $\sum_{\ell=1}^{L}\lambda_{\ell}=1$, and $f_{\ell}$’s are p-to-e calibrators. Here, p-to-e calibrator is a decreasing function $f:[0,\infty)\mapsto[0,\infty]$ such that $\|f\|_{L_{1}([0,1])}\leq 1$ (Vovk and Wang,, 2021; Gasparin et al.,, 2024). This form of p-value aggregation is proposed in Vovk et al., (2022). If $p_{\ell}$’s are conventional p-values, $G_{p}(\mathbf{p})=\inf\{\alpha\in(0,1):\mathbf{p}\in R_{\alpha}\}$ with $R_{\alpha}$ defined in (6) encompasses some popular aggregation methods. For instance, if $f_{\ell}(p)=2-2p$ and $\lambda_{\ell}=1/L$ for all $\ell\in[L]$, we get the arithmetic mean aggregation function $G_{p}(\mathbf{p})=2\sum_{\ell=1}^{L}p_{\ell}/L$; for a pre-determined $k\in[L]$, if $f_{\ell}(p)=L/k\cdot\operatorname{\mathds{1}}\{p\in(0,k/L)\}+\infty\operatorname{\mathds{1}}(p=0)$ and $\lambda_{\ell}=1/L$ for all $\ell\in[L]$, we obtain the Rüger’s method $G_{p}(\mathbf{p})=L/k\cdot p_{(k)}$.
Analogous to Proposition 3, the following proposition establishes the super-uniformity of the aggregated synthetic p-values.
Proposition 5.
Suppose (1) holds. Let $\bar{p}(Y):=G_{p}\{\mathbf{p}(Y)\}=\inf\{\alpha\in(0,1):\mathbf{p}(Y)\in R_{\alpha}\}$, where $R_{\alpha}$ is defined in (6). Then, we have $\mathbb{P}\{\bar{p}(Y)\leq t\}\leq t$ for all $t\in[0,1]$.
For synthetic e-values under dependence, aggregation can be achieved through convex combination. The theoretical guarantee is provided in the next proposition.
Proposition 6.
Suppose (1) holds. Let $\bar{e}(Y):=\bar{e}(Y;\lambda)=\sum_{\ell=1}^{L}\lambda_{\ell}e_{\ell}(Y)$, where $\lambda_{\ell}\geq 0$ for all $\ell$ and $\sum_{\ell=1}^{L}\lambda_{\ell}=1$. Then, we have $\mathbb{E}\{\bar{e}(Y)\}\leq 1$.
2.3 Test Inversion
As a final step, we transform the aggregated synthetic statistics back into an uncertainty set. It is well known that there is a correspondence between hypothesis testing and set estimation. In fact, a standard method for constructing confidence sets is through test inversion (Casella and Berger,, 2024). We state this classical result in the following proposition.
Proposition 7.
Suppose we have a $\alpha$-level test for the null hypothesis $H_{0}(\theta):\theta^{*}=\theta$ for each $\theta\in\Theta$. If we define $\mathcal{C}_{\alpha}=\{\theta:H_{0}(\theta)\ \text{is not rejected}\}$, then $\mathcal{C}_{\alpha}$ is a valid $1-\alpha$ confidence set.
The inversion principle similarly applies to constructing general uncertainty sets, as formally stated later in Proposition 8. Based on this observation, we perform the test inversion step of Algorithm 1 based on the aggregated p-values or e-values. Specifically, for any pre-specified $\alpha\in(0,1)$, the final merged set is constructed via
$\bar{\mathcal{C}}_{\alpha}=\{y:\bar{p}(y)>\alpha\}\text{ or }\bar{\mathcal{C}}_{\alpha}=\{y:\bar{e}(y)<\tau/\alpha\},$
where $\bar{p}(y)$ and $\bar{e}(y)$ are the aggregated statistics studied in Section 2.2.
Remark 4.
Consider the scenario where $\alpha_{\ell}=\alpha$ for all $\ell\in[L]$. If $\tau=1/2$ and $\bar{e}(y)=\frac{1}{L}\sum_{\ell=1}^{L}e_{\ell}(y)$, then the merged set $\bar{\mathcal{C}}_{\alpha}=\{y:\bar{e}(y)<\tau/\alpha\}$ reduces to the majority voting set $\bar{\mathcal{C}}_{\alpha}^{\tt MV}=\big{\{}y\in\mathcal{Y}:\frac{1}{L}\sum_{\ell=1}^{L}\operatorname{\mathds{1}}(y\in\mathcal{C}_{\ell,\alpha_{\ell}})\geq\frac{1}{2}\big{\}}$ proposed in Gasparin and Ramdas, (2024).
Remark 5.
Randomization techniques for power improvement, similar to those presented in Gasparin et al., (2024); Gasparin and Ramdas, (2024), can also be applied to our construction of synthetic statistics. Specifically, we can replace (6) with
$R^{\rm U}_{\alpha}=\left\{\mathbf{p}\in[0,1]^{L}:\sum_{\ell=1}^{L}\lambda_{\ell}\cdot f_{\ell}\left(\frac{p_{\ell}}{\alpha}\right)\geq U\right\},$
and $\bar{e}(Y;\lambda)$ with $\bar{e}(Y;\lambda)/U$, where $U$ is an independently generated $\rm{Unif}(0,1)$ random variable. By employing such randomization, we may obtain smaller merged sets with valid coverage rates. The validity of these procedures follows from Theorem 3.10 in Gasparin et al., (2024) and Theorem 1.2 in Ramdas and Manole, (2023).
2.4 SAT Procedure for Infinite Candidate Space
In many cases the candidate space $\mathcal{Y}$ is infinite or even uncountable, which makes it impractical to compute $\{p_{\ell}(y)\}_{y\in\mathcal{Y}}$ individually as specified in Algorithm 1. To address this, we propose a modified version of SAT in Algorithm 2. Specifically, we split $\mathcal{Y}$ into non-overlapping subsets and select a representative candidate from each subset to execute Algorithm 1. Note that Algorithm 2 is guaranteed to terminate in finite time if $L<\infty$ and $\mathcal{C}_{\ell,\alpha_{\ell}}$ are finite unions of connected sets.
Algorithm 2 The SAT Procedure for Practical Implementation
0:  The pairs $\{(\mathcal{C}_{\ell,\alpha_{\ell}},\alpha_{\ell})\}_{\ell\in[L]}$, candidate space $\mathcal{Y}$, a suitable aggregation function $G_{p}(\cdot)$ (or $G_{e}(\cdot)$), a control level $\alpha\in(0,1)$, an adjustment factor $\tau\in(0,1]$ (optional, take $\tau$=1 as default).
0:  $\bar{\mathcal{C}}_{\alpha}\leftarrow\{\}$, $\mathcal{M}_{0}\leftarrow\{\mathcal{Y}\}$.
1:  for each $\ell$ in [L] do
2:     Iteratively split $\mathcal{Y}$ into =$\mathcal{M}_{\ell}=\cap\{\mathcal{M}_{\ell-1},\mathcal{C}_{\ell,\alpha_{\ell}}\}\cup\setminus\{\mathcal{M}_{\ell-1},\mathcal{C}_{\ell,\alpha_{\ell}}\}$ for all $\ell\in[L]$, where $\cap\{\mathcal{A},b\}=\{a\cap b:a\in\mathcal{A}\}$ and $\setminus\{\mathcal{A},b\}=\{a\setminus b:a\in\mathcal{A}\}$.
3:  end for
4:  for each $\tilde{\mathcal{Y}}\in\mathcal{M}_{L}$ do
5:     Select any representative candidate $y\in\tilde{\mathcal{Y}}$.
6:     for each study $\ell\in[L]$ do
7:        Generate synthetic p-values $p_{\ell}(y)$ using (2) (or synthetic e-values $e_{\ell}(y)$ using (3)).
8:     end for
9:     Calculate $\bar{p}(y)=G_{p}\{\mathbf{p}(y)\}$ where $\mathbf{p}(y)=\{p_{1}(y),\ldots,p_{L}(y)\}$
                (or $\bar{e}(y)=G_{e}\{\mathbf{e}(y)\}$ where $\mathbf{e}(y)=\{e_{1}(y),\ldots,e_{L}(y)\}$).
10:     Update $\bar{\mathcal{C}}_{\alpha}\leftarrow\bar{\mathcal{C}}_{\alpha}\cup\tilde{\mathcal{Y}}$ if $\bar{p}(y)>\alpha$ (or if $\bar{e}(y)<\tau/\alpha$).
11:  end for
11:  Merged set $\bar{\mathcal{C}}_{\alpha}$.
3 Theoretical Analysis
In this section, we first establish the validity of the proposed SAT procedure. We then study the size of the merged set and the admissibility of deterministic SAT.
3.1 Validity of SAT
We begin with the inversion principle for general uncertainty sets, followed by a theorem that confirms the finite sample theoretical coverage guarantee of Algorithm 1. The validity of Algorithm 2 is established in Section A.1 of the supplement.
Proposition 8.
The following two statements hold.
  1. 1.
    For any $\alpha\in(0,1)$, define $\mathcal{C}_{\alpha}=\{y\in\mathcal{Y}:p(y)>\alpha\}$. If $\mathbb{P}\{p(Y)\leq t\}\leq t$ for all $t\in[0,1]$, then $\mathcal{C}_{\alpha}$ is a $(1-\alpha)$-level uncertainty set for $Y$.
  2. 2.
    For any $\alpha\in(0,1)$, define $\mathcal{C}_{\alpha}=\{y\in\mathcal{Y}:e(y)<\tau/\alpha\}$. If $\mathbb{E}\{e(Y)\}\leq 1$, then $\mathcal{C}_{\alpha}$ is a $(1-\alpha/\tau)$-level uncertainty set for $Y$.
The above result, together with Propositions 3 - 6, implies that Algorithm 1 produces a merged uncertainty set that achieves the target coverage rate.
Theorem 1.
If any of Propositions 3 - 6 holds and the corresponding aggregation function is applied, then the merged set $\bar{\mathcal{C}}_{\alpha}$ produced by Algorithm 1 (with the default choice of $\tau=1$) satisfies $\mathbb{P}(Y\in\bar{\mathcal{C}}_{\alpha})\geq 1-\alpha$ for any $\alpha\in(0,1)$.
3.2 Size of the Merged Set
We now turn to the expected size of the merged set produced by Algorithm 1; the theoretical results also apply to Algorithm 2, based on the extended definition of synthetic p-values or e-values as discussed in Section A.1 of the Supplement. Note that the expected size of $\bar{\mathcal{C}}_{\alpha}$ can be written as
$\mathbb{E}\left\{\int_{\mathcal{Y}}\operatorname{\mathds{1}}(y\in\bar{\mathcal{C}}_{\alpha})\,\mu({\rm d}y)\right\}=\int_{\mathcal{Y}}\mathbb{P}(y\in\bar{\mathcal{C}}_{\alpha})\,\mu({\rm d}y),$ (7)
where $\mu$ is an appropriate measure. For example, if $\mathcal{Y}=\mathbb{R}^{n}$, $\mu$ can be the Lebesgue measure; if $\mathcal{Y}$ is a discrete set, $\mu$ can be the counting measure. According to (7), analyzing $\mathbb{P}(y\in\bar{\mathcal{C}}_{\alpha})$ allows us to study the average size of $\bar{\mathcal{C}}_{\alpha}$. Note that, the set $\bar{\mathcal{C}}_{\alpha}$ is significantly influenced by the aggregation method used for combining synthetic statistics. We focus on the independent aggregation methods while the analyses of dependent scenarios are relegated to Section A.2 of the supplement. We start with synthetic p-value aggregation.
Assumption 1.
Assume that $S_{\ell}=S$ for all $\ell\in[L]$ and $S:[0,1]\mapsto\mathbb{R}$ satisfies:
  • (i)
    $\|S\|_{L^{2}([0,1])}\leq C_{S}$ for some constant $C_{S}>0$.
  • (ii)
    $\alpha^{-1}\int_{0}^{\alpha}S(t)\,{\rm d}t>(1-\alpha)^{-1}\int_{\alpha}^{1}S(t)\,{\rm d}t$ for all $\alpha\in(0,1)$.
We remark that Part $(ii)$ of Assumption 1 is automatically satisfied if $S$ is strictly decreasing. Assumption 1 is met by common choices of $S(t)$, such as $S(t)=-2\log(t)$ and $S(t)=-\Phi^{-1}(t)$. We establish the results on $\mathbb{P}(y\in\bar{\mathcal{C}}_{\alpha})$ in the following theorem.
Theorem 2.
Suppose (1) holds with $\alpha_{\ell}=\alpha$ for all $\ell\in[L]$, and $\mathcal{C}_{\ell,\alpha}$’s are independent and identically distributed. Let $\bar{p}(y)=\inf\{\alpha\in(0,1):\mathbf{p}(y)\in R_{\alpha}\}$ for each $y\in\mathcal{Y}$, where $R_{\alpha}$ is defined in (4). If $S_{\ell}$ satisfies Assumption 1 and $\operatorname{{\rm Var}}[S\{p_{1}(y)\}]>0$ for all $y\in\mathcal{Y}$, then for any $\alpha^{\prime}\in(0,1)$, when $L\rightarrow\infty$, there exists constant $C>0$ such that
$\mathbb{P}(y\notin\bar{\mathcal{C}}_{\alpha^{\prime}})=1-O\{\exp(-C\cdot L)\},\quad\forall y\in\{y\in\mathcal{Y}:\mathbb{P}(y\notin\mathcal{C}_{1,\alpha})>\alpha\},$
where $\bar{\mathcal{C}}_{\alpha^{\prime}}=\{y\in\mathcal{Y}:\bar{p}(y)>\alpha^{\prime}\}$.
To interpret Theorem 2, we consider the following simple example. Suppose $Y=\theta^{*}$ is a fixed unknown parameter, each of the $L$ studies independently draws a sample $X_{\ell}$ from $\mathcal{N}(\theta^{*},1)$ and constructs an uncertainty set $\mathcal{C}_{\ell,\alpha}=\{\theta\in\mathbb{R}:|\theta-X_{\ell}|\leq z_{1-\alpha/2}\}$. In this case, the set $\{\theta\in\mathbb{R}:\mathbb{P}(\theta\notin\mathcal{C}_{1,\alpha})>\alpha\}$ is simply \$\mathbb{R}\backslash\{\theta^{*}\}$ and Theorem 2 implies that for all $\theta\neq\theta^{*}$, the probability that $\bar{\mathcal{C}}_{\alpha^{\prime}}$ includes $\theta$ converges to 0 at a rate of $\exp(-C\cdot L)$ for any $\alpha^{\prime}\in(0,1)$. Consequently, the final merged set will converge to the singleton $\{\theta^{*}\}$ if $L\rightarrow\infty$ and becomes infinitesimal in size.
We next present the parallel result for synthetic e-value aggregation.
Theorem 3.
Suppose (1) holds with $\alpha_{\ell}=\alpha$ for all $\ell\in[L]$, and $\mathcal{C}_{\ell,\alpha}$’s are independent and identically distributed. Let $\bar{e}_{k}(y)=G_{e}\{\mathbf{e}(y);k\}$ be defined as in (5) for each $y\in\mathcal{Y}$. Then, for any fixed constant $k\in[L]$ and any $\alpha^{\prime}\in(0,1)$, when $L\to\infty$, there exists constant $C>0$ such that
$\mathbb{P}(y\notin\bar{\mathcal{C}}_{\alpha^{\prime}})=1-O\{\exp(-C\cdot L)\},\quad\forall y\in\left\{y\in\mathcal{Y}:\mathbb{P}(y\notin\mathcal{C}_{1,\alpha})>\alpha\left(\frac{\tau}{\alpha^{\prime}}\right)^{1/k}\right\},$
where $\bar{\mathcal{C}}_{\alpha^{\prime}}=\{y\in\mathcal{Y}:\bar{e}_{k}(y)<\tau/\alpha^{\prime}\}$ for any fixed $\tau\in(0,1]$.
For the default choice of $\tau=1$, the set $\{y\in\mathcal{Y}:\mathbb{P}(y\notin\mathcal{C}_{1,\alpha})>\alpha(\tau/\alpha^{\prime})^{1/k}\}$ is increasing in $k$, and is always a subset of $\{y\in\mathcal{Y}:\mathbb{P}(y\notin\mathcal{C}_{1,\alpha})>\alpha\}$ in Theorem 2 for any fixed $k$. This observation aligns with the established conservativeness of e-value aggregation (Barber et al.,, 2019; Blier-Wong and Wang,, 2024).
3.3 Admissibility of SAT
In the previous subsection, we examined the properties of both randomized and deterministic set merging procedures, and in particular, analyzed the size of the merged set. However, it remains an open question whether the SAT procedure can be strictly improved. This naturally leads to the concept of admissibility, which provides a formal criterion for optimality. In this section, we focus on the admissibility of deterministic set merging functions, where such theoretical guarantees are most meaningfully defined and can be rigorously established.
Definition 1.
A function $f$ is called a level-$\alpha$ deterministic set merging function if it takes a collection of $L$ uncertainty sets $\{\mathcal{C}_{1,\alpha_{1}},,\ldots,\mathcal{C}_{L,\alpha_{L}}\}$ and outputs a single set $\mathcal{C}_{\alpha}$ that satisfies $\mathbb{P}(Y\notin\mathcal{C}_{\alpha})\leq\alpha$, without introducing any external randomness. A level-$\alpha$ deterministic set merging function $f$ is called admissble if there does not exists another level-$\alpha$ deterministic set merging function $g\neq f$ such that $g(\cdot)\subseteq f(\cdot)$ for any valid inputs.
The following theorem establishes the admissibility of SAT.
Theorem 4.
Without additional assumptions, every admissible level-$\alpha$ deterministic set merging function can be represented in the form of SAT, utilizing synthetic e-values defined in (3) and a convex combination in the aggregation step.
Remark 6.
The converse of Theorem 4 is not true. In Remark D.1 of the Supplement we give a counterexample where SAT with synthetic e-values and certain choice of convex combination is not admissible.
The main idea in proving Theorem 4 is to exploit the duality between uncertainty sets and the e-function defined in (3). This duality allows us to recast the problem of merging uncertainty sets as that of merging e-functions—a powerful perspective that enables us to draw on established results from the e-value literature for deeper insight. A key step in the proof of Theorem 4 is to show that all admissible e-function mergers that map collections of e-functions to a single e-function must take the form of convex combinations. We emphasize that this result is not a direct consequence of Theorem 1 in Wang, (2025), as our focus here is on merging functions of a specific form rather than general random variables. In fact, our proof of Theorem 4 uses different techniques from those in the existing e-value literature.
In the special case where all $L$ initial uncertainty sets share the same miscoverage level, it is natural to consider symmetric set merging functions—those that are invariant under permutations of the input sets. The following theorem establishes the admissibility of SAT under this setting.
Theorem 5.
Under the assumption that all $L$ initial uncertainty sets have the same miscoverage level, the SAT procedure with synthetic e-values and arithmetic mean aggregation yields the only admissible symmetric deterministic set merging function.
4 Simulation Studies
In this section, we study the empirical performance of the proposed methods on simulated datasets. Throughout the paper, we abbreviate the methods using the format ($\cdot$)+($\cdot$). The first component denotes the type of synthetic statistic: SyP and SyE correspond to the synthetic p-value and e-value, respectively. The variant SyP(naïve) refers to a direct construction of the synthetic p-value via $\alpha_{\ell}\cdot\operatorname{\mathds{1}}(y\notin\mathcal{C}_{\ell,\alpha_{\ell}})+\operatorname{\mathds{1}}(y\in\mathcal{C}_{\ell,\alpha_{\ell}})$, as discussed in Remark 2. Additionally, OrP denotes the “oracle p-value,” which is used to generate the initial uncertainty sets. The second part refers to the aggregation methods, which we summarize in Table 1. In the last step of inverting the synthetic e-values, we set $\tau=1$. When all initial uncertainty sets are independent of each other and have the same coverage level, we include the procedure in Section 2.6 of Gasparin and Ramdas, (2024) for comparison, which is denoted as MV Binom.
Method Abbreviation Aggregation function
P-value Fisher’s method Fisher $1-F_{\chi^{2},{2L}}\big{(}-2\sum_{\ell=1}^{L}\log p_{\ell}\big{)}$
Arithmetic mean AM $2\sum_{\ell=1}^{L}p_{\ell}/L$
Rüger’s method with $k=1$ Rüger ${L}\cdot p_{(1)}\cdot\operatorname{\mathds{1}}(p_{(1)}>0)$
E-value Arithmetic Mean AM $\sum_{\ell=1}^{L}e_{\ell}/L$
Equation (5) with $k=2$ U2 $\binom{L}{2}^{-1}\sum_{\mathcal{I}_{2}\in\mathcal{B}_{2}}\prod_{\ell\in\mathcal{I}_{2}}e_{\ell}(Y)$
Table 1: A summary of aggregation methods.
The following four scenarios are considered.
  1. Scenario 1.
    $L=5$, $\alpha_{1}=\cdots=\alpha_{L}=\alpha/2$ varies from $0.01$ to $0.1$;
  2. Scenario 2.
    $L$ varies from $2$ to $9$, $\alpha_{1}=\cdots=\alpha_{L}=\alpha/2=0.05$;
  3. Scenario 3.
    $L=5$, $(\alpha_{1},\ldots,\alpha_{L})=(0.01,\ldots,0.05)$, $\alpha$ varies from $0.01$ to $0.1$;
  4. Scenario 4.
    $L=5$, $\alpha_{1}=\cdots=\alpha_{L}$ varies from $0.01$ to $0.1$, $\alpha=0.1$.
For all of the settings, we simulate $L\in\mathbb{N}$ initial uncertainty sets $\mathcal{C}_{\ell,\alpha_{\ell}}$ for $\ell\in[L]$ with individual set coverage guarantee: $\mathbb{P}(Y\notin\mathcal{C}_{\ell,\alpha_{\ell}})\leq\alpha_{\ell}$. Our goal is to construct a merged set $\bar{\mathcal{C}}_{\alpha}$ that satisfies $\mathbb{P}(Y\notin\bar{\mathcal{C}}_{\alpha})\leq\alpha$. All experiments are based on 5000 replications, and the average results are reported.
4.1 Merging Independent Uncertainty Sets
We let $Y=2$ be a fixed parameter. For each of the $L$ studies, we independently draw $n=3$ samples from $\mathcal{N}(Y,1)$, and denote the mean of the $n$ samples obtained by study $\ell$ as $\bar{X}_{\ell}$. The oracle p-value at each candidate point $y$ is computed as $p^{\text{or}}(y)=2\Phi\left(-\sqrt{n}|y-\bar{X}_{\ell}|\right)$, and the corresponding uncertainty set $\mathcal{C}_{\ell,\alpha_{\ell}}$ is constructed by $\mathcal{C}_{\ell,\alpha_{\ell}}=\left[\bar{X}_{\ell}-z_{\alpha_{\ell}/2}/\sqrt{n},\bar{X}_{\ell}+z_{\alpha_{\ell}/2}/\sqrt{n}\right].$ The comparison results of various methods are summarized in Figure 2.
Figure 2: Coverage and size of the merged uncertainty sets for the normal mean estimation problem. Each individual set is constructed based on a two-sided $z$-test.
From the top panel of Figure 2, we observe that all methods successfully control the desired coverage level. The size comparisons presented in the second row indicate that SyP+Fisher demonstrates clear advantages over other variations of SAT, and the observed size trends align with the theoretical results in Theorem 2 and Theorem 3.
4.2 Merging Dependent Uncertainty Sets
Consider the linear model $Y=X^{\top}\beta+\epsilon$, where $X\in\mathbb{R}^{150}$ is the covariate vector, $\beta\in\mathbb{R}^{150}$ is an unknown vector of coefficients, and $\epsilon$ is the error term that follows $\mathcal{N}(0,1)$. In this experiment, we generate the first 10 entries of $\beta$ from $\mathcal{N}(0,4I_{10\times 10})$ and set the remaining entries to 0. We then generate 400 pairs of data according to the linear model, with $X$ sampled independently from $\mathcal{N}(0,I_{150\times 150})$. Next, we sample one more $X$ from $\mathcal{N}(0,I_{150\times 150})$, and our goal is to produce an uncertainty set for the corresponding $Y$. In this setting $Y$ is random, so the synthetic p-values and e-values are not independent. Consequently, aggregation methods like U2 and Fisher are no longer valid, and we do not include them for comparison.
Conformal prediction set with different learning algorithms
We consider the case where each study chooses different learning algorithms to construct conformal prediction sets. More precisely, for study $\ell$ the non-conformity score for a candidate $y$ is $|y-\hat{f}_{\ell}(X)|$, and $\hat{f}_{\ell}$ is one of the following models: neural network, random forest, LASSO and linear regression. 200 pairs of data are randomly picked as training data, and the rest are used as calibration data. All studies use the same split. We use the package conformalInference [^3] Code is provided in https://github.com/ryantibs/conformal. to implement these methods. Since there are only four learning algorithms, we have $L=4$ in this experiment and Scenario 2 is omitted. The result is summarized in Figure 3. Similar to Section 4.1, all variations of SAT achieve the target coverage rate for the merged set. We observe that SyP+Rüger slightly outperforms SyE+AM in some cases. Note that this does not contradict the theory presented in Section 3.3, as SyP+Rüger is a randomized merging procedure, whereas Theorems 4 and 5 pertain exclusively to deterministic merging procedures. In contrast, both SyP(naïve)+AM and SyP(naïve)+Rüger are deterministic merging procedures. As shown in Figure 3, the resulting sizes from these two procedures are consistently larger than those of SyE+AM, which numerically confirms our theoretical results. For more sophisticated deterministic p-value merging procedures, see Section B of the Supplement. The simulation outcomes in Figure B.1 further support the admissibility conclusion.
Figure 3: Coverage and size of the merged conformal prediction sets evaluated using different score functions and merging methods. The initial sets are constructed using a full conformal approach, with neural network, random forest, LASSO, and linear model selected as the score functions.
Conformal prediction set with different splits of training and calibration data
In this experiment, we merge split conformal prediction sets that are constructed using the same learning algorithm but different splits of the training and calibration data. The non-conformity for an candidate $y$ is $|y-\hat{f}_{\ell}(X)|$ and $\hat{f}_{\ell}$ is obtained from a LASSO model. Each study randomly selects 200 data points from the 400 labeled samples for training and uses the remaining 200 points for calibration, denoted as $D^{(\ell)}_{\rm tr}$ and $D^{(\ell)}_{\rm cal}$, respectively. The results are summarized in Figure 4. It shows that all variations of SAT successfully achieve the target coverage level, with SyE+AM performing comparably to SyP+Rüger, and both uniformly outperforming SyP(naïve)+AM and SyP(naïve)+Rüger.
Figure 4: Coverage and size of the merged conformal prediction sets by different splits of the training and calibration data. The initial sets are constructed using a split conformal approach with LASSO selected as the score function.
5 Real Data Analysis
[…]
Figure 5: Coverage and size of the merged prediction sets using different learning algorithms for dataset ImageNet__\__val.
The results are summarized in Figure 5. In comparison to the simulated data from the first part of Section 4.2, which considers similar setups in linear models, the image data and initial set constructions in the current real data context are significantly more complex. Nonetheless, since SAT relies solely on the generated initial sets, its overall performance is similar to that shown in Figure 3, demonstrating the effectiveness and robustness of the proposed procedures.
6 Discussions
In this paper, we introduced the SAT framework for merging uncertainty sets in settings where only the initial uncertainty sets and their corresponding control levels are available. The proposed method is flexible, computationally efficient, and requires minimal information from each individual study.
The size of the merged set produced by SAT critically depends on the aggregation method used to combine synthetic statistics. As shown in Theorem 5, the SAT procedure with synthetic e-values and arithmetic mean aggregation yields the only admissible symmetric deterministic set merging function. However, our simulation studies reveal that certain randomized symmetric merging procedures can, in some cases, outperform SAT with synthetic e-values and arithmetic mean aggregation. Exploring admissibility in the presence of randomization thus presents a compelling direction for future research.
Another important direction for future work is to investigate the admissibility of SAT under independence assumptions. We conjecture that when synthetic p-values are independent, SAT combined with Fisher’s method yields an admissible procedure. This hypothesis remains to be rigorously established and will require further theoretical explorations.
The SAT procedure begins by converting uncertainty sets into synthetic statistics. Similarly, it is feasible to convert multiple testing results into synthetic statistics. Consequently, a procedure similar to SAT could be employed to aggregate each study’s rejection set, effectively controlling false discoveries. This idea is briefly discussed in Section C.2 of the supplement, but a comprehensive understanding warrants further explorations.